OpenStudy (anonymous):

Write a rule for the linear function whose graph has a slope m = - 1/3, and a y-intercept b= -2 a. y = 1/3x -2 b. y = -1/3x - 2 c. y = -3x - 2 d. y = 1/3x + 2

Nnesha (nnesha):

what is \the slope intercept form equation ?

OpenStudy (anonymous):

y = mx +b

Nnesha (nnesha):

slope is what and y-intercept is what in that equation ?? :-)

OpenStudy (anonymous):

-1/3 +2 ... ?

OpenStudy (anonymous):

or -2

OpenStudy (anonymous):

ohhh hahaha

Nnesha (nnesha):

i'm asking in this equation y = mx+ b what is slope and what is y-intercept ?

OpenStudy (anonymous):

the y-intercept is -2

Nnesha (nnesha):

so m is slope and b is y-intercept right so just replace m and b by their values

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y= -1/3 -2 ?

Nnesha (nnesha):

yep right

Nnesha (nnesha):

wait just replace m by -1/3 there suppose to be x

OpenStudy (anonymous):

y = -1/3x -2?

Nnesha (nnesha):

yep!